Intercontinental’s special order requires 1,600 kilograms of genatope, a solid chemical regularly used in the company’s products. The current stock of genatope is 9,300 kilograms at a book value of 8.20 p per kilogram. If the special order is accepted, the firm will be forced to restock genatope earlier than expected, at a predicted cost of 8.80 p per kilogram. Without the special order, the purchasing manager predicts that the price will be 8.40 p when normal restocking takes place. Any order of genatope must be in the amount of 6,600 kilograms.

What is the relevant cost of genatope?
Answer #1

SOLUTION

Amount ($)
Replacement cost of genatope 13,440
Cost of rush order 2,640
Total relevant cost 16,080

Replacement cost of genatope = Quantity required * Price for normal restocking

= 1,600 * 8.40 = 13,440

Cost of rush order = Quantity ordered * (Rush order price - Normal price)

= 6,600 * (8.80 - 8.40)

= 6,600 * 0.4 = 2,640